### v1

Abstract
Large Language Models (LLMs) demonstrate strong reasoning and task planning capabilities but remain fundamentally limited in physical interaction modeling. Existing approaches integrate perception via Vision-Language Models (VLMs) or adaptive decision-making through Reinforcement Learning (RL), but they fail to capture dynamic object interactions or require task-specific training, limiting their real-world applicability. We introduce APEX (Anticipatory Physics-Enhanced Execution), a framework that equips LLMs with physics-driven foresight for real-time task planning. APEX constructs structured graphs to identify and model the most relevant dynamic interactions in the environment, providing LLMs with explicit physical state updates. Simultaneously, APEX provides low-latency forward simulations of physically feasible actions, allowing LLMs to select optimal strategies based on predictive outcomes rather than static observations. We evaluate APEX on three benchmarks designed to assess perception, prediction, and decision-making: (1) Physics Reasoning Benchmark, testing causal inference and object motion prediction; (2) Tetris, evaluating whether physics-informed prediction enhances decision-making performance in long-horizon planning tasks; (3) Dynamic Obstacle Avoidance, assessing the immediate integration of perception and action feasibility analysis. APEX significantly outperforms standard LLMs and VLM-based models, demonstrating the necessity of explicit physics reasoning for bridging the gap between language-based intelligence and real-world task execution. The source code and experiment setup are publicly available at https://github.com/hwj20/APEX_EXP .
1 Introduction
A cat is about to pounce on an LLM-controlled agent. The agent detects the cat nearby and knows it should move, but does it understand that the cat will jump in 2 seconds? Once the LLM decides to evade, multiple escape routes exist, how does it choose a path that avoids both the cat and surrounding obstacles? These two challenges: understanding dynamic interactions and predicting action consequences, highlight fundamental limitations in existing LLM-based agents.
Current methods attempt to address these issues using Vision-Language Models (VLMs)  (Wang et al., 2024a; Ahn et al., 2022; Huang et al., 2024, 2023; Liang et al., 2023; Liu et al., 2024; Hu et al., 2023b) and Reinforcement Learning (RL) (Patel et al., 2025; Lee et al., 2024; Ma et al., 2024a). However, they remain fundamentally limited:
  • Static Perception Without Dynamic Awareness: VLMs enable LLMs to recognize objects but fail to model interactions over time—they can detect a cat but cannot anticipate its movement. In real-world decision-making, static snapshots are insufficient; understanding object motion is essential.
  • Lack of Action-Outcome Feedback and Physical Grounding: Existing approaches often treat decision-making as a one-shot prediction task, offering no structured feedback loop between actions and their physical consequences. Instead of modeling the environment’s response through grounded physical equations, they rely on latent dynamics (World Models) or reward-driven adaptation (RL). As a result, these systems lack interpretable, quantitative feedback on action feasibility, e.g., whether an action would cause collision, balance failure, or violate timing constraints.
  • Expensive and Slow Policy Adaptation: RL-based approaches, such as VoxPoser (Huang et al., 2023) and Code-as-Policies (Liang et al., 2023), require extensive task-specific training. Every new scenario demands costly retraining, making real-time adaptation impractical.
Figure 1: Comparison of physical reasoning capabilities across three systems—LLM without spatial grounding, VLM and world modeling, and our proposed APEX—on three scenarios involving object prediction, agent-object interaction, and action planning. While vanilla LLMs are not necessarily making random choices in the prediction task, our experimental results in Section 4 indicate that their performance is statistically indistinguishable from random selection in this context. APEX provides not only qualitative predictions but also quantitative estimations of outcomes (e.g., time to impact, risk of collision), demonstrating its structured understanding of physical causality.
To plan actions in the real world, agents must do more than perceive and react—they must simulate, quantify, and foresee. We introduce APEX (Anticipatory Physics-Enhanced Execution), a framework that enables LLMs to anticipate environmental changes and optimize actions through physics-based reasoning. APEX constructs structured graphs that extract the most relevant dynamic interactions in an environment(Nishida et al., 2018; Huang et al., 2025), enabling LLMs to reason about object motion and forces. Additionally, APEX performs future state simulation(Smith et al., 2013), predicting how different actions will alter the environment over time, providing explicit physical constraints to guide decision-making. This strengthens the standard LLM’s capabilities in physical reasoning, perception, and prediction, empowering LLM-driven agents to perform low-latency planning in physical environments, as illustrated in Fig. 1.
We evaluate APEX across three benchmark tasks, each is designed to address a critical limitation in existing approaches:
  • Physics Reasoning Benchmark (Addressing Static Perception): Testing LLMs’ ability to infer object dynamics beyond simple object recognition.
  • Tetris (Evaluating Physics-Driven Foresight): Testing whether providing forward physical simulations as feedback improves the long-horizon decision quality of language models in structured planning environments.
  • Dynamic Obstacle Avoidance (Addressing Real-Time Adaptation): Assessing real-time integration of perception and prediction for adaptive decision-making, ensuring LLMs can dynamically adjust their behavior based on future state simulations.
We aim to close the gap between language-based reasoning and physically grounded execution. Our contributions are:
  • APEX: a unified framework that equips LLMs with real-time perception with graph networks and physical foresight for dynamic task planning.
  • A three-part benchmark suite spanning structured reasoning, long-horizon planning, and real-time control, each targeting a distinct dimension of physical intelligence.
  • Empirical results showing that APEX outperforms LLMs and VLM-based agents in (1) numerical reasoning and physical calculation (Physics QA); (2) simulation-guided planning with physical intuition (Tetris); and (3) perception-integrated prediction for real-time decision-making (dynamic obstacle avoidance).
2 Related Work
Despite significant progress in task planning for LLM-based or VLM-based agents (Wang et al., 2024b; Ma et al., 2024b; Kawaharazuka et al., 2024; Hu et al., 2023a), existing paradigms largely fail to integrate real-time physical modeling in embodied intelligence. Our work is situated at a unique intersection of language reasoning, graph-based physical abstraction, and online physics simulation.
2.1 Vision-Language Models: Perception Without Physical Consequence
Vision–language models (VLMs) such as CLIP (Radford et al., 2021), Flamingo (Alayrac et al., 2022), PaLM-E (Driess et al., 2023), and OpenVLM (Kim et al., 2024) learn powerful image–text embeddings that support zero-shot recognition and instruction following (Ma et al., 2024b).
Many VLM-empowered agents, such as CLIPort (Shridhar et al., 2022), VIMA (Jiang et al., 2022), VoxPoser (Huang et al., 2023), RT-2 (Brohan et al., 2023), and PhysVLM (Zhou et al., 2025)—inherit this same static worldview. They augment visual grounding with spatial transport layers, multimodal prompting, or feasibility masks, but still cannot generalize to novel dynamics and remain blind to explicit physical laws. A handful of works have tried to close the loop by training Transformer-based action predictors directly on VLM features, for example, RT-1 (Brohan et al., 2022) learns end-to-end vision-to-control policies. More recently, DeepMind’s generalist agent Gato (Reed et al., 2022) showed that a single Transformer can handle images, text, and control signals in a unified framework.
Yet these approaches still encode physics only implicitly in learned weights, offering no transparent physical feedback and often failing under distributional shifts.
2.2 World Models: Foresight Without Guarantees
Subsequent extensions further enrich this paradigm: TWM (Robine et al., 2023) incorporates temporal attention into latent rollouts; SMART (Sun et al., 2023) adds self-supervised multi-task pretraining for control; R3M (Nair et al., 2022) leverages universal visual representations; and Genie (Bruce et al., 2024) integrates interactive environment generation. These models introduce video representation learning (Majumdar et al., 2023), multi-agent dynamics, and forward/inverse prediction, yet all remain black-box latent estimators without explicit guarantees of physical consistency.
Despite their imaginative capabilities, latent world models exhibit fundamental limitations, including compounding roll-out errors that exacerbate over extended horizons, poor robustness to distributional shifts, and opaque latent dynamics that obscure failure modes and hinder interpretability. Furthermore, their temporal abstraction often distorts real physical intervals by embedding time into latent structures rather than modeling it explicitly.
2.3 Reinforcement Learning: Expensive Mastery, Poor Generalization
Reinforcement learning (RL) algorithms such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) and Soft Actor–Critic (SAC) (Haarnoja et al., 2018), and Imitation Learning (IL) like Generative Adversarial Imitation Learning (GAIL) (Ho and Ermon, 2016) have achieved notable success in robotic control through extensive trial-and-error interaction. More recent LLM‐guided RL hybrids aim to mitigate these issues by combining language reasoning with policy learning. SayCan (Ahn et al., 2022) uses a language model to rank actions proposed by a pretrained policy, and Inner Monologue (Huang et al., 2022) adds on‐the‐fly replanning via chain-of-thought prompts. Iker (Patel et al., 2025) augments low‐level controllers with iterative keypoint rewards from a VLM.
Models such as RT-1 Brohan et al. (2022) and BC-Z Jang et al. (2022) demonstrate the potential of large Transformer policies to generalize across multiple tasks after extensive pretraining on diverse environments. ProgPrompt (Singh et al., 2023), PromptCap (Hu et al., 2023c), and ECOt (Zawalski et al., 2024), chain LLM reasoning for task planning.
Despite these advancements, RL and its LLM‐centric extensions still face substantial challenges. They either require millions of environment steps to converge, struggle under physical distribution shifts, or rely on predefined controllers with limited adaptability to new physics interactions.
2.4 Physical Simulation in LLM Reasoning: Beyond Conceptual Hallucinations
Prior works like Mind’s Eye (Liu et al., 2022) and PiLoT (Zhang et al., 2023) propose injecting simulation-derived hints into LLM prompts to correct conceptual hallucinations, such as misunderstandings of qualitative physics (e.g., “heavier objects fall faster”). While effective for symbolic reasoning, these methods overlook a key dimension: numerical precision. Our experiments show that modern LLMs (e.g., GPT-4o(Achiam et al., 2023)) already grasp qualitative physical rules, limiting the value of such corrections. However, they still fail at quantitative tasks—like predicting collision timing—unless grounded by external computation. In real-world environments where timing and magnitude are critical, this gap is consequential.
Table 1 lists the comparison and positioning between current methods and our framework.
Table 1: Comparison of Planning Approaches in Dynamic Physical Environments
Method Perception Dynamic Awareness Prediction Quant. Physics Real-Time Generability
Vanilla LLMs
VLMs
World Models
RL
APEX (Ours)
3 Methodology
Figure 2: Overview of the APEX reasoning pipeline. Visual input is abstracted into a motion-aware interaction graph via DG-Motion Attention. This graph structure triggers simulation in a physical engine (MuJoCo), which evaluates the outcome of candidate actions. A vanilla LLM then selects an action based on the simulated consequences. This loop, graph → trigger → simulate → LLM → act, enables grounded, temporally, aware physical reasoning.
In this section, we introduce the APEX framework, structured explicitly in five detailed stages as shown in Fig. 2, systematically integrating physical reasoning into LLM decision-making. At its core, APEX leverages a graph-based representation explicitly chosen for its inherent ability to model relationships, not merely to highlight the most immediate or obvious actions but rather to capture complex, task-relevant interactions comprehensively.
3.1 Graph: Relational Scene Representation
Given consecutive snapshots at times $t$ and $t+\Delta t$, we construct relational graphs $G_{t}$ and $G_{t+\Delta t}$ over the same set of object nodes. Each node corresponds to a distinct entity in the environment, and edges encode potential interactions between pairs of objects. This relational graph structure explicitly represents the complex web of interactions, emphasizing task-relevant relationships rather than isolated physical states.
Such graph representations can be directly connected to upstream 3D reconstruction modules, serving as an intermediate abstraction layer between raw perceptual input and structured physical reasoning.
3.2 Trigger: Difference-Graph Motion Attention
We form a difference graph:
$\Delta G=G_{t+\Delta t}-G_{t},$
whose edges encode per-pair displacement, relative velocity, and newly emerging or evolving relationships. A Graphormer encoder computes attention scores:
$\alpha_{ij}=\mathrm{Graphormer}(G_{t},G_{t+\Delta t})_{ij},$
identifying the most task-relevant edges based on relational dynamics. The selected edges define a focused subgraph, which is translated into a concise natural-language summary $S$, explicitly describing critical interactions and relationships (e.g., "sphere A is about to collide with B, influencing agent strategy").
3.3 Simulate: Physics-Grounded Action Rollouts
From the current relational state $s_{t}$, we enumerate a discrete set of candidate actions $\{a_{i}\}$ (e.g., left, right, down, jump). Note that while we define the action set $A$ as a collection of potential actions, the size of $A$ remains limited due to the finite degrees of freedom in current robotic systems, making enumeration feasible (Glover, 2004; Sutton et al., 1998). For each candidate action $a_{i}$, we invoke forward simulations:
$s_{t+1}^{(i)}=\mathrm{PhysicsSim}(s_{t},a_{i}),$
and generate outcome descriptors $r_{i}$ (collision flags, target distances, object positions, durations). These outcome descriptors offer explicit, physics-grounded feedback tied directly to relational predictions and task implications.
3.4 LLM: Guided Decision Synthesis
We enrich the original LLM prompt $x$ with the relational summary $S$ and detailed simulation outcomes $\{r_{1},\dots,r_{n}\}$, resulting in a contextually comprehensive prompt:
$x^{\prime}=x\cup S\cup\{r_{i}\}.$
The augmented context guides the LLM to synthesize the optimal relationally-informed decision sequence $\Pi$, representing a series of actions strategically selected to achieve the target objective based on predictive outcomes:
$\Pi^{\prime}=\arg\max_{\Pi}P_{\mathrm{LLM}}(\Pi\mid x^{\prime}).$
3.5 Act: Execution of the Optimal Plan
The action plan $\Pi^{\prime}$ is executed in the environment, realizing robust interactions grounded explicitly in relationally-informed physical foresight.
Algorithm 1 APEX: Anticipatory Physics-Enhanced Execution
0: Environment snapshots at $t$ and $t+\Delta t$, LLM prompt $x$
0: Final LLM-generated action plan $\Pi^{\prime}$
1: Construct relational graphs $G_{t}=(V,E)$ and $G_{t+\Delta t}$ from object states
2: Compute attention scores via Graphormer:
$\alpha_{ij}=\text{Graphormer}(G_{t},G_{t+\Delta t})_{ij}$
3: Identify top-$k$ relationally salient edges forming focused subgraph $\tilde{G}$
4: Generate summary $S$ from relational interactions within $\tilde{G}$
5: Enumerate feasible actions $\{a_{1},\dots,a_{n}\}$ from current relational state
6:for each action $a_{i}$ do
7:  Simulate future state: $s_{t+1}^{(i)}=\text{PhysicsSim}(s_{t},a_{i})$
8:  Generate outcome description $r_{i}=\text{Describe}(s_{t+1}^{(i)})$
9:end for
10: Append summary $S$ and outcomes $\{r_{1},\dots,r_{n}\}$ to LLM prompt, forming enriched prompt $x^{\prime}$
11: Decode optimal action plan from LLM:
$\Pi^{\prime}=\arg\max_{\Pi}P_{\text{LLM}}(\Pi\mid x^{\prime})$
12:return $\Pi^{\prime}$
4 Experiments
To evaluate APEX, we introduce a new LLM Physical Reasoning Benchmark, testing AI models’ ability to predict and adapt to dynamic environments. The evaluation consists of three primary experiments as shown in Table 2.
Table 2: Summary of Experimental Setups and Physical Reasoning Capabilities
Experiment Capability Verified Evaluation Objective
Physics Reasoning Physical reasoning over multiple entities Test LLM’s ability to understand motion-related quantities across targets.
Tetris Planning Foresight via simulated prediction Assess whether physics-informed feedback improves planning quality.
Obstacle Avoidance Perception-integrated prediction Validate perception-action grounding under dynamic environments.
4.1 Experiment 1: Physical Reasoning Accuracy in Structured Tasks
To assess the physical reasoning capabilities of LLMs, we construct a suite of synthetic 3D tasks grounded in classical mechanics, including linear motion, circular motion, projectile motion, multi-object interactions, and collision prediction. Each task is framed as a structured reasoning problem: given object positions, velocities, and physical parameters, the LLM must infer whether collisions will occur or predict resulting velocities after interaction.
We compare vanilla GPT-4o against our APEX-enhanced GPT-4o and report three metrics:
  • Accuracy: Whether the model provides a fully correct structured answer within the tolerance of 5%.
  • Mean Squared Error (MSE): Quantitative deviation from ground-truth numerical values.
  • Numerical Validity: Percentage of fields where the model returns valid numbers.
Table 3 summarizes the results. GPT-4o struggles in multi-body and circular dynamics due to limited temporal abstraction and lack of grounded physical priors. In contrast, APEX achieves near-perfect accuracy and precision, demonstrating that explicit modeling significantly boosts physical reasoning performance.
We conduct ablation experiments on different $dt$ in the physical simulation engine with the Euler forward method on Table 4. Here, $dt$ refers to the step size in the physics engine’s forward simulation, not the time interval in the Graph Trigger module.
Table 3: Comparison of GPT-4o vs APEX-enhanced GPT on Physical Reasoning Tasks
Task Type Accuracy (%) $\uparrow$ MSE $\downarrow$ Numerical Validity (%) $\uparrow$
GPT-4o
3D Linear Motion 8.00 213.5931 28.00
3D Circular Motion 24.00 4.0998 76.00
3D Projectile Motion 88.00 303.6022 100.00
3D Collision 44.00 12.4816 100.00
3D Multi-Object Motion 0.00 1918.2065 81.33
APEX (ours)
3D Linear Motion 96.00 0.0076 100.00
3D Circular Motion 100.00 0.0000 100.00
3D Projectile Motion 100.00 0.0001 100.00
3D Collision 88.00 2.4627 100.00
3D Multi-Object Motion 97.33 0.0013 100.00
Table 4: Simulation accuracy and average duration per question type at different timesteps dtdt.
[…]
3D Linear Motion 100.00 0.023 100.00 0.0058 96.00 0.0042
3D Circular Motion 100.00 0.028 100.00 0.0068 96.00 0.0046
3D Projectile Motion 92.00 0.013 92.00 0.0035 48.00 0.0027
3D Multi-Object Motion 97.33 0.076 90.67 0.022 80.00 0.013
3D Collision 98.00 0.0073 98.00 0.0090 98.00 0.0080
4.2 Experiment 2: Real-time Physical Planning in Tetris
We design a second benchmark to test the agent’s ability to perform dynamic, physics-informed planning in a classic block-stacking domain: Tetris. Unlike traditional planning tasks that focus on symbolic correctness or visual alignment, this environment emphasizes physical feasibility, spatial reasoning, and long-horizon optimization.
The agent interacts with a Tetris simulator in which it must select actions (left, right, rotate, drop) for falling blocks. Each decision must be made based on the current board state, the shape of the block, and the anticipated physical consequences of different placements. All models are run under the same sequence of five randomized seeds, and each episode is capped at 15 blocks with the estimated maximum number of clear lines as 3, ensuring fair and bounded comparison.
We compare different decision systems:
  • GPT-4o & GPT-4o-mini: baseline LLMs with no physical modeling.
  • VLM: GPT-4o with images as the VLM (Patel et al., 2025; Wang et al., 2025) that perceives the current board state via screenshot input.
  • APEX (Ours): physical planning with physics-based rollout.
We evaluate each model on five physically grounded metrics:
  • Final Score: total score after game termination (each cleared line counts 100).
  • Max Height: the tallest column reached during gameplay.
  • Hole Count: number of empty cells beneath landed blocks.
  • Bumpiness: total height difference between adjacent columns.
  • Height Increase per Move: average vertical growth rate per action.
These metrics reflect task performance and physical efficiency jointly. A low bumpiness and hole count indicate stable and compact stacking, while a lower height delta per move demonstrates the agent’s foresight in minimizing vertical sprawl.
Table 5: Comparison of Baselines vs APEX on Tetris-style Structured Planning
Model Final Score $\uparrow$ Max Height $\downarrow$ Holes $\downarrow$ Bumps $\downarrow$ Height $\Delta$/Move $\downarrow$
GPT-4o 0.0 14.6 33.4 25.6 +14.6
GPT-4o-mini 0.0 18.2 26.0 36.4 +18.2
VLM 0.0 12.6 30.2 22.6 +12.6
APEX (ours) 140.0 5.0 2.8 6.8 +5.0
4.3 Experiment 3: Dynamic Obstacle Avoidance
This experiment assesses the agent’s adaptive decision-making capabilities within dynamic physical environments characterized by moving obstacles. The setup utilizes a simulated MuJoCo environment where an LLM-driven agent navigates through varying obstacle densities and speeds across different difficulty levels.
The evaluation metrics are as follows:
  • CFR (Collision-Free Rate): the percentage of time in which the agent successfully avoids all obstacles.
  • IAR (Invalid Action Rate): the proportion of actions that lead to collisions or unsafe states.
  • AST (Average Survival Time): the average duration the agent remains operational without colliding, reflecting overall navigation efficacy.
Table 6: Performance on Real-time Obstacle Avoidance Across Task Complexities
Model Simple Medium Hard
[…]
GPT-4o-mini 0 0 2.55 0 0 1.86 0 0 2.24
GPT-4o 20 0 5.85 0 0 3.15 0 0 1.66
VLM 0 7 5.18 0 4 3.14 0 7 2.48
APEX (GPT-4o-mini) 100 0 10.00 60 0 8.64 20 0 6.86
APEX (GPT-4o) 100 0 10.00 100 0 10.00 60 0 8.07
We conduct ablation experiments on different graph models and different values of k on Table 7.
Table 7: Ablation Study on Hard Obstacle Avoidance: Top-$k$ Selection vs. Graph Model Choice
[…]
Top-$k$ $k=1$ gpt-4o-mini 0 7.17 0 0.74
$k=2$ gpt-4o-mini 0 6.38 0 0.73
$k=4$ gpt-4o-mini 0 7.53 0 0.74
$k=1$ gpt-4o 0 4.97 0 0.94
$k=2$ gpt-4o 40 6.58 0 1.25
$k=4$ gpt-4o 40 5.58 0 1.29
Graph Model GAT gpt-4o-mini 0 2.85 0 0.89
GCN gpt-4o-mini 0 7.08 0 0.85
Graphormer gpt-4o-mini 0 7.59 0 0.76
GAT gpt-4o 0 4.99 0 0.62
GCN gpt-4o 0 5.26 0 0.70
Graphormer gpt-4o 20 5.44 0 1.24
4.4 Evaluation Summary
APEX substantially augments LLM capabilities in physical reasoning across structured tasks, dynamic adaptation, and real-time obstacle avoidance. Our findings indicate that APEX consistently outperforms standard LLMs, achieving over 90% accuracy in multi-object dynamics (Experiment 1), efficient long-horizon planning (Experiment 2), and proactive collision avoidance (Experiment 3).
In Experiment 1, APEX demonstrates superior accuracy in predicting circular motion and collision dynamics, with baseline GPT-4o achieving less than 20% on Table 3.
In Experiment 2 (Tetris), APEX leverages predictive foresight to minimize structural irregularities, optimizing placements and significantly improving task performance on Table 5.
Experiment 3 further underscores APEX’s advantage in real-time obstacle avoidance, effectively mitigating collision risks through predictive modeling, a critical gap in baseline GPT and VLM systems on Table 6.
5 Conclusion
In this paper, we introduced APEX, a novel framework that enhances LLMs with predictive physical reasoning capabilities by integrating graph-based physical modeling, and physics simulation. Unlike prior methods that rely on static observations or constraint filtering, APEX enables LLMs to anticipate future physical interactions and adapt task plans accordingly. Experimental results demonstrate that APEX significantly improves performance on physical reasoning benchmarks, outperforming standard LLMs, VLM-based task planning, and grounded decoding techniques.
Furthermore, APEX’s structured approach to physical modeling opens new opportunities for future research in AI-driven task planning, robotics, and autonomous decision-making. This study provides a new perspective on enhancing LLMs’ physical reasoning capabilities by replacing RL-based trial-and-error learning with predictive physical modeling. This direction presents new possibilities for future robotic task planning and can be combined with existing VLM+RL-based methods to further improve LLMs’ ability to handle physical interaction tasks.
6 Future Work
As a next step, we aim to extend APEX into a unified world modeling framework, APEX++, where the language model serves not only as a planner, but as a core component in a recurrent, interpretable perception-prediction-action loop. This would allow for the emergence of grounded intelligence capable of proactive behavior, structured foresight, and physical adaptability, unlocking new possibilities across robotics, self-driving, and embodied AI.
Ultimately, we envision APEX++ as a general-purpose cognitive architecture, blending symbolic reasoning with physical grounding, to guide intelligent agents through complex, dynamic environments with a shared understanding of how the world works.